# Cloud Storage Integrity at Scale:

### v2

@@ VII-B. Results
    How well can state-of-the-art hash tree designs scale with capacity, cache size, and read/write ratio?
@@
    To what extent can DHTs or state-of-the-art designs deliver optimal performance?
@@
    What are the memory and storage overheads associated with implementing DHTs?
Refer to captionRefer to caption
(capacity=1TB, rratio=1%, I/O size=4KB)     (capacity=1TB, cache=10%, I/O size=4KB)    (capacity=1TB, cache=10%, rratio=1%)
Refer to caption Refer to captionRefer to caption
 (capacity=1TB, rratio=1%, I/O size=4KB)     (capacity=1TB, cache=10%, I/O size=4KB)    (capacity=1TB, cache=10%, rratio=1%)
Figure 8. Write (top) and read (bottom) performance under the zipf2.5 workload w.r.t. cache size, read ratio, and I/O size.
Scaling with Capacity. Figure 7 shows the IOPS and latencies observed across various capacities and under two workloads. The Zipfian experiment is configured with a read ratio of 1%, which reflects the read ratio observed in the Alibaba trace. We make three key observations.
First, both read and write performance of binary and 64-ary hash trees decreases logarithmically with capacity. Under the Zipfian workload, they observe up to 3.5$\times$ and 4$\times$ IOPS slowdowns over the no encryption/no integrity baseline, respectively. Under the Alibaba workload, they also incur increasing slowdowns as capacity scales, with up to a 40.1$\times$ slowdown at 8TB capacity. The average latencies also show the same pattern. Notably, 64-ary trees, which are state-of-the-art for secure memory systems, incur up to a 6.3$\times$ slowdown under the Zipfian workload and a 35.9$\times$ slowdown under the Alibaba workload. Note that 64-ary trees must be initialized with a capacity that is a power of 64. We also note that since the traces are scaled to our experiment capacity, IOPS generally decreases (due to scaled I/O sizes being larger) but the effective bandwidths remain the same.
In contrast, DHTs provide up to a 2.3$\times$ and 4.2$\times$ write IOPS speedup over binary and 64-ary hash trees, respectively. They also provide up to 2.7$\times$ and 4$\times$ write latency speedups over binary trees and 64-ary trees, respectively. Importantly, even as capacity increases, DHTs can reliably deliver the same amount of IOPS. These results show that static nature of state-of-the-art methods cannot capture patterns in workloads, and DHTs can more reliably scale to larger capacities. In fact, DHTs deliver ¿70% of the baseline throughput across all studied capacities under the Zipfian workload. Although they also provide ¿1.7$\times$ speedups over the state-of-the-art under the Alibaba workload, overheads compared to the insecure baseline are still high in general, due to I/O sizes being larger. Future work could extend balanced trees and DHTs to support (safe) delaying/batching of writes to efficiently support large write requests. However, as reflected in the Zipfian workload, when workloads are mostly composed of small 4KB writes (which occurs most often in practice (Li et al., 2023)), DHTs deliver the lowest overhead compared to the insecure baseline.
We now focus on the Zipfian traces to more broadly explore the parameter space.
Scaling with Cache Size. Figure 8 (left) shows performance with respect to cache size. Note that we specify cache size as a ratio of the tree size, and therefore the absolute cache size varies with the storage capacity. We make one key observation: small caches are very efficient. As the cache size increases beyond 1%, throughputs observed under each hash tree design do not increase. Yet, at a cache size ratio of 100% (which is unrealistic in practice, but is useful for modeling a perfect cache), DHTs still observe the same speedups. Thus, all hash tree designs make efficient use of available cache memory, but the structure of DHTs enables them to provide higher performance at a lower cache budget (we discuss this in more detail below). This shows that DHTs can provide large speedups at various operating points for cache size.
Scaling with Read/Write Ratio. Figure 8 (middle) shows how the performance changes with respect to the read/write ratio. We expect that at higher read ratios, DHTs, balanced, and optimal trees will all observe higher absolute throughputs, as reads can be quickly served by early returns due to caching. However, we see a very sharp performance drop for balanced trees when there is a significant proportion of writes. Note that writes are expensive becauase, unlike read requests, writes must always be serialized to the root to ensure consistency with the root hash and the data disk. At a read ratio of 50%, we see that writes become costly for balanced trees: DHTs are 1.7$\times$ faster.
At a read ratio of 1%, DHTs are 2.3$\times$ faster. The latter case is most often observed in practice, largely because higher-level caches (e.g., application caches) can efficiently serve reads, leaving I/Os sent to storage to be mostly writes. Notably, even when workloads are read-heavy, DHTs deliver comparable performance to the state-of-the-art. We attribute this to the fact that DHTs inherit the theoretical guarantees of splay trees, which ensure $O(logn)$ amortized lookup (i.e., verification or update) time. Thus, DHTs can perform at least as good as the balanced trees on average, for either read- or write-heavy workloads.
Scaling with I/O Size. Figure 8 (right) shows how performance changes with respect to I/O size. With increased I/O size, IOPS generally decreases, but bandwidth remains the same. We make two key observations. First, DHTs provide ¿2.3$\times$ speedups over balanced trees across each I/O size. With large 128K I/Os, the slowdowns of each hash tree design with respect to the insecure baseline increases due to increased per-I/O costs. In the context of the insecure baseline, the block layer in the kernel can mitigate per-I/O costs through optimizations like write buffering/batching. Balanced hash trees and DHTs could both be extended with (safe) write buffering/batching to reduce the gap with the insecure baseline. However, we note that the speedups observed by DHTs should scale proportionally.
Achieving Optimal Performance. Our optimal model helped informed the design of DHTs but also allows us to measure a concrete upper bound on performance. Figure 7 and Figure 8 also characterize the performance of balanced trees and DHTs with respect to the optimal. Note that we provide optimal results only for the Zipfian traces. The process of scaling the Alibaba traces (e.g., scaling a 1TB disk trace down to a trace for a 1GB disk) introduces a loss in fidelity in the traces and affects the Huffman algorithm in unpredictable ways. However, we note that constructing the tree under the original trace yields a tree that provides maximum throughput. For example, constructing our optimal hash tree over a 500GB disk from the original trace of volume #4 shown in Figure 7 (which was recorded on a 500GB disk) yields a tree that provides 23K IOPS.
With respect to capacity, we observe for the Zipfian workload that DHTs deliver ¿95% of optimal performance (¿42K IOPS) as capacity scales up, while balanced trees provide only 55% of optimal performance in the best case at 1GB. With respect to cache size, read ratio, and I/O size, we observe similarly that DHTs deliver ¿95% optimal throughput across all experiments, while the slowdown factor between balanced trees and the optimal increases with capacity, write-heaviness, and I/O size. This shows that balanced hash trees do not provide a reliable guarantee on performance, while DHTs provide a stable performance guarantee. We attribute this to DHTs being based on splay-trees, which are conjectured to provide performance within a constant factor of the optimal tree structure under any access pattern.
@@
TABLE II. DHTs require additional memory/storage to cache/store each tree node. However, DHTs break even on this trade-off, because they can provide higher performance than balanced trees, at a smaller cache budget.
@@
Memory & Storage Overhead. While DHTs provide several advantages, they have higher memory and storage requirements than balanced trees. Table II shows the memory overheads for DHTs. DHTs cannot use implicit indexing like balanced trees, but instead require explicitly storing parent-child pointers (as integer node IDs) both for nodes in-memory and on-disk. This implies at least one additional integer field for leaf nodes, at most three additional integer fields for internal nodes, and one additional integer hotness counter field for all nodes.
This may have adverse affects on cache performance, since less nodes in a DHT can be stored within the same cache budget as a balanced tree. However, we show that cache hit rates are very high even for very small caches. Moreover, DHTs provide better performance at a smaller cache budget than balanced trees at a higher cache budget. For example, at a cache budget of $x$ for balanced trees, DHTs would occupy $(1.44x+1.8x)/2=1.62x$ cache memory (and storage), but if DHTs provide a speedup $>1.62x$, then we can conclude that, in terms of monetary cost, DHTs provide better performance per bit of cache memory purchased and thus make more efficient use of available cache memory. As we have shown, DHTs provide up to 4.2$\times$ speedups over balanced trees.
Key takeaway: Our experiments show that under a realistic set of system configurations and storage workloads, overheads incurred by state-of-the-art hash trees are prohibitive. Meeting the demands of large-scale, high-throughput, low-latency cloud storage requires a fundamental redesign of integrity mechanisms. Our observations of optimal tree structures and DHTs sheds light on how we can break the scalability limit of prior solutions.
VIII. Security Analysis
Below we analyze how DHTs remain resilient against the integrity violations outlined in Section III.
Data Correctness. Like traditional balanced hash tree designs (Taassori et al., 2018), DHTs authenticate block data with the MACs that serve as the leaves in the hash tree. The MACs ensure the correctness of data, and are checked whenever data is read or written by clients. Violating correctness guarantees would require a successful MAC forgery attack, which is infeasible under standard cryptographic assumptions.
Spatial Uniqueness. In DHTs, encryption and MACs are keyed to the block address by using the address as additional authenticated data (AAD). This ensures that data contents have not been corrupted, and also that the data is associated with the address intended by the client (i.e., blocks are not swapped before being returned to the client).
Temporal Uniqueness. Like balanced hash trees (Figure 2), DHTs prevent replay attacks at the block layer by authenticating all reads and writes against the root hash. The root hash is the trust anchor of the system. While the trees are structurally different, the root hash serves the same purpose.
IX. Discussion
Below we discuss additional considerations regarding DHTs.
Leaf Order Consistency. The most significant difference from traditional balanced trees is that, in DHTs, nodes may be rotated to arbitrary locations in the tree. In some domains, such as blockchains, the leaf order carries meaning. For example, transactions represent leaves and must be organized in some well-defined order. DHTs would violate the integrity of such an ordering and would therefore not be suitable for such a context. In the context of cloud block storage, order does not matter. That is, nodes can be placed arbitrarily in the tree, insofar as the hash tree invariants outlined in Section VI are met. Consistency in leaf ordering is thus a domain-specific property. For storage, the hash tree structure itself needs only provide the conduit to authenticate reads and writes against the root hash. As nodes are moved to different locations in the tree, we still ensure spatial uniqueness by keying encryption and MAC generation to unique block addresses.
Block Size. Like other works, our basic unit of storage is a typical 4KB block. Some recent works have explored using larger block sizes to improve the performance of storage systems (e.g., 1 MB blocks for object storage). Evidently, using larger blocks can improve bandwidth utilization. However, there is no one-size-fits-all solution when choosing a suitable block size. For example, with larger blocks, if the application’s I/O size does not align with the block size, then the system may suffer from I/O amplification. In the Alibaba dataset, we also found that a large majority I/Os (¿90%) are small (4KB) for most volumes. We leave future work to exploring this trade-off in more detail.
X. Related Work
@@
Hash trees have become a core component of many computing systems, including blockchains, secure memories, etc. (Buterin, 2016; Naor and Nissim, 2000; Android, 2023; Taassori et al., 2018; Avanzi et al., 2022; Gassend et al., 2003; McKeen et al., 2013; Erway et al., 2015; Arasu et al., 2021; Sinha and Christodorescu, 2018; Li et al., 2006). We discuss these related works below.
@@
Secure Memory Systems. Secure memories have been long-studied (Yan et al., 2006; Gassend et al., 2003; Feng et al., 2021; Rogers et al., 2007). Their central goal is to provide a secure environment in which the secrecy and integrity of application code and data can be assured. Recently, secure memories have seen widespread commercial success through implementations such as Intel SGX (McKeen et al., 2013; che Tsai et al., 2017; Priebe et al., 2020). At the heart of a secure memory system is the hash tree, which ensures all three security properties discussed in Section III. It is known that hash trees can severely impact performance, and optimizing them has thus been a central focus of recent research.
State-of-the-art approaches, such as Penglai (Feng et al., 2021), FastVer (Arasu et al., 2021), and VAULT (Taassori et al., 2018), have shown that various optimizations like caching, increasing tree fanout, and parallelization, can ensure low-overheads in some contexts (e.g., at small capacities). However, there are several key differences between these works and our work. First, we focus on persistent storage rather than main memory, which is subject to vastly different workload characteristics, capacities, and assumptions about cache memory size. We considered the state-of-the-art for secure memories (high-degree trees(Freij et al., 2021; Rogers et al., 2007)) for juxtaposition, and show that such approaches are not applicable to storage. Notably, with larger fanouts, the number of auxiliary hashes that need to be fetched during an update or verification increases proportional to the fanout, thus requiring additional hash fetches (and reauthentications on misses). Moreover, some prior works have decidedly side-stepped the issue of writes by either suppressing caches during updates or limiting analysis to read-heavy or read-only workloads (Arasu et al., 2021; Sinha and Christodorescu, 2018). Yet, storage workloads are known to be write-heavy, and we therefore closely examine write-heavy workloads to understand their implications on scalability.
DHTs take a fundamentally different approach to optimizing the hash tree: instead of increasing tree fanout, we remove the balance constraint to reduce the height of authentication paths. Specifically, high-degree trees trade-off increased overall fetch costs for reduced height, wheres DHTs trade-off increased cold-access fetch costs for reduced height.
Further, experiments in prior works have also been limited to simulation-based analysis and against small capacities (e.g., 16GB). We report experiments with an implementation on real block devices and consider a wide range of system settings: up to several TB capacity, with various cache sizes, with various read/write ratios. Other works using techniques such as lazy verification have sacrificed security guarantees for performance (Arasu et al., 2021). Because we focus on a more practical threat model, we do not compare against these works. However, we note that DHTs could be extended with such techniques, but leave such an analysis to future work.
@@
Authenticated Data Structures. Hash trees have also been examined in the broader theoretical context of authenticated data structures—data structures that have strong integrity protections (Tamassia, 2003; Miller et al., 2014)—for a variety of real-world problems. They become a central component of mobile and embedded device storage: dm-verity has played a pivotal role in providing verified boot for Android smartphones (Android, 2023). Other works have also examined hash trees in the context of blockchains (Buterin, 2016), certificate revocation systems (Melara et al., 2015), and provable data possession schemes (Erway et al., 2015; Dahlberg et al., 2016).
These works have highlighted the theoretical efficiency of different approaches to ensuring integrity of arbitrary datasets, but there is consensus between the theory and systems communities that Merkle hash trees are state-of-the-art method (Crosby and Wallach, 2011). Our work builds on these efforts by showing that, while efficient in theory, traditional static, balanced hash trees still incur substantial overheads in real systems, and this motivates the design of DHTs.
XI. Conclusion
We showed that state-of-the-art hash tree designs incur significant performance overheads and cannot scale to meet the demands of cloud storage systems. We addressed this problem by introducing a definition of an optimal hash tree, an oracle methodology to measure whether a hash tree design performs optimally, and a novel hash tree design called DHTs that exploits characteristics of optimal trees to provide up to a 4.2×4.2$\times$4.2 × speedup on read and write performance over the state-of-the-art. Our contributions provide a new foundation in the search for integrity mechanisms that can operate efficiently at scale. Notably, we re-characterize the integrity problem from a static one into a dynamic one, showing the viability of integrity structures that learn and exploit patterns in workloads to improve performance. Our code is plug-and-play into standard Linux systems and is open-sourced at [Anonymized link].

### v3

@@ VII-B. Results
    How well do state-of-the-art hash tree designs perform across the various system and workload settings that characterize cloud block storage deployments?
@@
    To what extent can DMTs improve performance over the state of the art, and under what conditions?
@@
    What memory and storage trade-offs do DMTs make?
Figure 11. Aggregate throughput with Read ratio at 1%. DMTs provide up to $\times2$ higher throughput than the state-of-the-art, demonstrating its ability to scale well with capacity.
Figure 12. P50 (top) and P99.9 (bottom) write latency. DMTs median and tail latencies reflect throughput improvements, demonstrating it can provide a stable performance guarantee.
Scaling with Capacity. We first analyze how disk capacity (which affects tree height) and hash cache size affect performance. Where appropriate, default parameters include—Read ratio: 1%, I/O size: 32 KB, Thread count: 1, I/O depth: 32, Capacity: 64 GB, Cache size: 10%. We choose these parameters because they showcase the best performing configuration for the baselines. We examine various workload shapes ranging from uniform to highly skewed Zipfian. We focus particularly on $\theta$: 2.5 because it closely approximates the shape of real-world storage workload patterns (li2023depth, ; yang2016write, ) (see Figure 18).
Figure 11 shows that aggregate read/write throughput decreases w.r.t. capacity for all balanced trees. Note that the Zipfian workload is emitted from an i.i.d. source and is therefore an exact upper bound. We observe that 64-ary trees are the worst performing: reduced tree height can reduce the effective number of hashes that must be computed, but results in lower cache efficiency, which amplifies metadata I/O costs. The state-of-the-art binary trees incur up to a 75% throughput loss over the Encryption/no integrity baseline at 4 TB. 4-ary and 8-ary trees similarly suffer from a 70% throughput loss at 4 TB. In contrast, DMTs consistently deliver the highest throughput and ¿85% of optimal throughput across all capacities. This clearly demonstrates that DMTs can scale to higher or lower capacities more efficiently. And as noted in Section IV, with even faster storage devices, the proportion of time spent hashing vs. performing the data access will grow substantially, increasing our observed DMT speedups.
Latency improvements are the same—DMTs splay on 1% of accesses and those costs are amortized over time because splaying occurs most frequently on hot data. Figure 12 corroborates this: DMT median and tail write latencies are still significantly lower than the state of the art.
Figure 13. Aggregate read/write throughput. DMTs perform best under skewed workloads; under uniform workloads they observe a 6% cost over binary trees due to exploratory splays.
Figure 14. Aggregate throughput. DMTs maintain the highest throughputs across both small and large cache sizes.
Impact of Workload Skewness. Figure 13 shows how performance changes w.r.t. workload skewness. We observe again that 64-ary trees are the worst performing. DMTs provide up to 2×2$\times$2 × speedups over the state of the art binary trees under heavy skew, but incur a 6% cost under more uniform patterns due to exploratory splays which yield no benefit. We attribute this low cost to the fact that DMTs inherit the theoretical guarantees of splay trees, which provide $O(log~{}n)$ amortized lookup (i.e., verification or update) time. Thus, DMTs perform at least as good as balanced (binary) trees on average.
We also observe that 4-ary and 8-ary trees deliver 25% higher throughput than DMTs under more uniform workloads. As discussed in Section IV, low-degree trees hit the optimal points in the design space for balanced trees (reduced tree height without adverse effects on cache performance); prior works have not considered this. However, when workloads become skewed, there is a substantial opportunity cost: 4-ary and 8-ary trees do not perform better than the optimal binary tree. This highlights a key observation: increasing tree degree alone is not sufficient to maximize performance. We believe that extending the DMT design to 4-ary and 8-ary trees will yield the most performant and generalized solution.
Figure 15. DMTs show speedups across different read ratios, I/O sizes, thread counts, and I/O depths.
Impact of Cache Size. The above analysis showed that DMTs can exploit skewed patterns in workloads when present and deliver a stable performance guarantee across various capacities. Now we examine the effects that other system settings and workload characteristics have on performance. The goal is to evaluate whether these conclusions about DMTs hold broadly. We continue with the Zipf(2.5) workload.
Figure 14 shows that DMTs maintain the highest throughputs across both small and large hash caches. Note that cache size is specified as a ratio of the tree size; the absolute cache size varies with capacity. Further, caches mostly benefit read I/Os (because they enable early returns), but write I/Os still must traverse the entire path to the root. In general, we observe that increasing cache size beyond 0.1% does not yield significant performance improvements for any hash tree design; small caches are already very efficient. Yet, losses observed by all balanced tree designs are still significant. This shows that caching only helps to an extent—when caches are efficient, hash tree overheads are largely attributable to how efficient the tree structure is. However, DMTs still deliver near-optimal performance and the highest across all sizes.
Impact of Read Ratio, I/O Size, Thread Count, and I/O Depth. Figure 15 (top) shows how the performance changes with respect to the read ratio. We expect that at higher read ratios, DMTs, balanced, and optimal trees will all observe higher absolute throughputs, as reads can be quickly served by early returns due to caching. However, when there is a significant proportion of writes ($\leq 50\%$ read ratio), DMTs provide nearly 2×2$\times$2 × higher throughput than balanced trees. Since storage access patterns tend to be write-heavy (due to application-level caches and the OS page cache), this shows that DMTs can more reliably handle write-heavy workloads, while delivering comparable performance under read-heavy workloads.
Figure 16. DMTs can adapt quickly to changing workloads, exploiting skewed patterns when present.
The remaining graphs in Figure 15 reflect the above observations. Baseline throughputs increase w.r.t. I/O size, but hash tree performance saturates at 32 KB I/Os—larger I/O sizes only lead to increased latencies without improved throughputs. A single thread is sufficient to saturate the device bandwidth. And an application I/O depth of 32 is sufficient to saturate the device bandwidth. DMTs still deliver up to 2×2$\times$2 × and 4×4$\times$4 × higher throughput over the state of the art binary and 64-ary trees. The extent to which DMTs see speedups is thus best attributed to the workload shape (Figure 13).
As noted in Section IV, 32 KB write I/Os require 8 sequential hash tree updates. State of the art works are not truly concurrent, but still rely on a global tree lock to serialize tree updates (taassori2018vault, ; freij2021bonsai, ; feng2021scalable, ). Performance has been shown to improve under high concurrency by using lazy verification (deferring and batching updates) (arasu2021fastver, ), but lazy verification violates freshness guarantees. Designing concurrency-optimal hash trees (and search trees in general (shanny2022occualizer, )) is an open problem.
Figure 17. On an Alibaba cloud volume trace, DMTs deliver notable speedups, while high-degree trees perform worst.
DMT dm-verity No enc/no integrity
write 255.4 MB/s 151.9 MB/s 318.8 MB/s
read 0.7 MB/s 0.4 MB/s 1.0 MB/s
TABLE II. Application read/write throughputs for the Filebench OLTP workload. DMT driver-level improvements are reflected at application-level.
Handling Changing Access Patterns. We now demonstrate that DMTs self-adapting nature is robust to changing workload patterns. Figure 16 shows a 150-second snapshot of sampled throughputs under a workload that exercises an extreme case where patterns alternate between uniform and skewed: Zipf(2.5) ¿ Uniform ¿ Zipf(2.0) ¿ Uniform ¿ Zipf(3.0). Phases are 30 seconds long, and the Zipfian phases are randomly centered at a new region in the address space. We observe that DMT throughput spikes within a few seconds of entering the Zipfian phases and DMTs maintain the speedup throughout. This shows that DMTs can capitalize on skewed patterns very quickly to maximize performance while delivering performance comparable to binary trees otherwise. As noted, extending DMTs to 4-ary trees can help further improve DMT performance during uniform phases.
Case study: Alibaba Cloud Volumes. We showed that DMTs perform near-optimally under a broad range of system and workload settings. We now examine how these observations hold under a real workload sampled from a recently published Alibaba trace dataset (li2023depth, ) (logical volume ID 4). We scale the offsets and I/O sizes proportionally to the experiment capacity. Note that the remaining volume traces are qualitatively the same (mean write ratio ¿98% and highly skewed). Further, note that the workload is non-i.i.d. and therefore H-OPT can underestimate the upper bound on throughput; temporal patterns enable DMTs to perform better in some cases.
Figure 17 (left) shows the aggregate throughputs observed at a 4 TB capacity, and Figure 17 (right) shows the distribution of write throughputs (sampled at 1-second intervals). Binary trees observe a 75% throughput loss at 4 TB capacity, while 64-ary trees observe an 88% throughput loss. DMTs provide a 1.3×1.3$\times$1.3 × speedup over the binary trees and a 1.2×1.2$\times$1.2 × speedup over the 4-ary trees. Importantly, the optimal (binary) tree still observes 15% higher throughput than 4-ary and 8-ary trees. This further supports our claim that a balanced tree structure is not sufficient to maximize the performance potential. As noted above, we believe the extending DMT principles to a 4-ary tree can help achieve maximum performance.
Case study: OLTP Workload. We now consider the Filebench OLTP workload, an exemplar application that is commonly run in the cloud and requires robust security protections (dong2024cloud, ). The goal is to evaluate how DMT device-level improvements translate to application-level improvements. The workload consists of 10 writer threads and 200 reader threads and is write-heavy. We run the workload for 10 minutes on a 1 TB disk (with a dataset size of $\approx$922 GB) formatted with ext4 and using a hash cache size of 10%. Table II shows that DMTs have 1.8×1.8$\times$1.8 × improved read and 1.7×1.7$\times$1.7 × improved write performance over the state of the art.
@@
TABLE III. DMTs require additional memory/storage for tree nodes, but break even on this trade-off: they provide higher performance than balanced trees, at a smaller cache budget.
@@
Memory & Storage Overhead. DMTs provide several advantages, but have higher memory and storage requirements than balanced trees (Table III). DMTs cannot use implicit indexing like balanced trees, but instead require explicitly storing parent-child pointers (as integer node IDs) both for nodes in-memory and on-disk. This implies at least one additional integer field for leaf nodes, at most three additional integer fields for internal nodes, and one additional integer hotness counter field for all nodes. However, we showed that cache hit rates are very high even for very small caches. For example, DMTs provide better performance at a cache size of 0.1% than binary trees do at a cache size of 1%. Thus, DMTs deliver better performance per dollar spent on cache memory.
Key takeaways: State-of-the-art hash trees incur substantial performance loss (up to 80%). DMTs address this by exploiting skewed patterns when present and adapting quickly to changes over time. We conclude that balanced trees are ill-suited as the base construction of a hash tree: maximizing performance requires tailoring the tree structure to workload patterns.
VIII. Related Work
@@
Hash trees are a core component of many computing systems, including blockchains, secure memories, etc. (buterin2016ethereum, ; naor2000certificate, ; android-dm-verity, ; taassori2018vault, ; avanzi2022cryptographic, ; gassend2003caches, ; mckeen_innovative_2013, ; erway2015dynamic, ; arasu2021fastver, ; sinha2018veritasdb, ; li2006dynamic, ; feng2021scalable, ). We discuss these related works below.
@@
Secure Memory. Secure memories have been long-studied (yan2006improving, ; gassend2003caches, ; feng2021scalable, ; rogers2007using, ). Their goal is to provide a secure environment in which the secrecy and integrity of application code and data can be assured. Recently, secure memories have seen widespread commercial success through implementations such as Intel SGX (mckeen_innovative_2013, ; tsai_graphene-sgx_nodate, ; priebe_sgx-lkl_2020, ). Consequently, recent work has shown that hash trees can severely degrade memory performance, and optimizing them has been a central focus of recent research. State-of-the-art approaches, such as Penglai (feng2021scalable, ), FastVer (arasu2021fastver, ), and VAULT (taassori2018vault, ), have shown that optimizations like caching and increasing tree degree can lower overheads in some contexts (e.g., at small capacities).
However, we focus on persistent storage rather than main memory. Hash tree performance implications in the context of cloud block storage remain largely unknown, perhaps due to the presumption that hashing costs are negligible at the block layer. We showed that this is not the case. Further, storage systems are subject to vastly different workload characteristics, capacities, and cache behaviors than memory systems. We juxtaposed DMTs against the high-degree trees used for secure memories (freij2021bonsai, ; taassori2018vault, )) and showed that such approaches are not applicable to storage. Additionally, some prior works have decidedly side-stepped the issue of writes by either suppressing caches during updates or limiting analysis to read-heavy or read-only workloads (arasu2021fastver, ; sinha2018veritasdb, ). Given that storage workloads are write-heavy, we closely examine write-heavy workloads. Further, experiments in these prior works have been limited to small capacities (e.g., 16GB). We report experiments with an implementation on real block devices on a live Linux system and evaluate a wider range of system settings.
@@
Authenticated Data Structures. Hash trees have also been examined in the broader theoretical context of authenticated data structures (tamassia2003authenticated, ; miller2014authenticated, ; crosby2011authenticated, ). They become a central component of mobile and embedded device storage: dm-verity has played a pivotal role in providing verified boot for Android smartphones (android-dm-verity, ). They have also been examined in the context of blockchains (buterin2016ethereum, ), certificate revocation systems (melara2015coniks, ), and provable data possession schemes (erway2015dynamic, ; dahlberg2016efficient, ). These works have highlighted the theoretical efficiency of Merkle hash trees. Our work shows that, while efficient in theory, traditional hash trees still incur substantial overheads in real systems. This motivates our search for a more efficient tree structure and ultimately the design of DMTs.
IX. Conclusion
Figure 18. Workload distributions.
Merkle hash trees provide robust integrity protections over untrusted storage, but they can severely degrade performance. We performed a comprehensive analysis of performance overheads, demonstrated the root cause, and designed an optimized tree structure called DMTs that advance the state of the art. DMTs demonstrate the power of integrity structures that exploit workload patterns to reduce overheads. Our code is open-sourced at https://github.com/MadSP-McDaniel/dmt.
Acknowledgments
This work was supported in part by the Semiconductor Research Corporation (SRC) and DARPA.